# Efficient4D: Fast Dynamic 3D Object Generation from a Single-view Video

### v4

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. An intuitive approach is to extend previous image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling. However, this approach would be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model. Importantly, our method can achieve real-time rendering under continuous camera trajectories. To enable robust reconstruction under sparse views, we introduce inconsistency-aware confidence-weighted loss design, along with a lightly weighted score distillation loss. Extensive experiments on both synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D. Our code and video results are publicly available at https://fudan-zvg.github.io/Efficient4D/.
keywords:
4D Generation, Efficiency, Diffusion Model, Gaussian Splatting, Video
@@ 1 Introduction
Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time. Despite the significant advancement of 3D object generation, existing works [1, 2, 3, 4] mostly consider static scenes or objects. With the availability of large-scale 3D datasets [5, 6], training generalizable models capable of directly generating multi-view images becomes possible [7, 8, 9]. These generated images can be turned into a 3D object through reconstruction techniques [10, 11]. By further augmenting these generated static objects with predefined animations [12], dynamic 3D content can be generated. However, this approach is still limited due to the need for fine-grained meshes as well as rigid restrictions.
@@
Directly generating 4D object/scene content from text description has been recently attempted [13]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [10] via score distillation sampling [1] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [14] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [13] this method is also slow to train (e.g., 120 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
@@
To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [8]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a 4D Gaussian splatting model [15]. This is an extension of the 3D Gaussian splatting [16], originally designed for static 3D scene representation, with the temporal dimension introduced additionally, allowing for real-time rendering under continuous camera trajectories. Using Gaussian representation brings about further computational efficiency gain, when compared with NeRF based designs (Figure 1). To tackle the challenging discontinuity between the generated sparse frames, we design an inconsistency-aware loss function based on whether there is confidence in the consistency of a frame with its adjacent frames. Along with a lightly weighted score distillation sampling loss for smooth viewpoint transitions, this technique enables robust 4D reconstruction. Notably, although our method takes a single-view video as input, Efficient4Dcan be easily extended to image-to-4D task by leveraging a image-to-video diffusion model [17].
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. We also extends Efficient4D to image-to-4D task. (iii) We introduce a inconsistency-aware confidence-weighted loss for reconstructing a 4D Gaussian splatting model using the generated training data efficiently and robustly. (iv) Extensive experiments on both synthetic and real data validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis. Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
Input Consistent4D [14]: 120 mins Efficient4D (Ours): 10 mins
Input 4DGen [18]: 130 mins Efficient4D (Ours): 10 mins
Input STAG4D [19]: 70 mins Efficient4D (Ours): 10 mins
@@ 2 Related work
3D generation
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [1], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [2, 20, 21, 22, 3, 23, 24, 25, 26, 4, 27] and image-to-3D tasks [28, 29, 30, 31] across various dimensions. However, SDS-based approaches often suffer from difficulty to converge and extended optimization times. Conversely, some efficient methods [32, 33, 7, 34, 8] have emerged. Notably, Point-E [32] and Shap-E [33] train models to directly generate 3D point clouds or meshes. Zero123 [7] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [34]. Importantly, SyncDreamer [8] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
[…]
4D representation
Efforts to synthesize videos with free-viewpoint control in dynamic scenes have a well-documented history [35]. For example, pre-NeRF [10] approaches challenges in reconstructing intricate scene details. Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [36], DeVRF [37] and HyperNeRF [38], which decouple geometry and motion, utilizing a canonical space and a learned deformation field. DynIBaR [39] deploys an image-based rendering paradigm for representing long videos with complex camera and object motions. Another group of methods [40, 41, 42] adopt tensor decomposition of 4D volumes to represent dynamic 3D scenes.
Recently, Gaussian Splatting [16] has received widespread attention for its real-time high-fidelity rendering, especially its explicit point-based representation which holds great potential in modeling dynamic scenes. Consequently, a significant amount of work has been proposed to explore its extension to dynamic scene modeling. Among them, Deformable 3D Gaussians [43] and 4DGaussian [44] integrated the deformation field with 3D Gaussian Splatting for the joint learning of the scene’s geometry and dynamics. SC-GS [45] represents the motion with a set of sparse control points to achieve reconstruction and motion editing. Unlike the previous radiance field-based representations often involving complex training schedules or suffering from slow convergence, these methods can be optimized efficiently, achieving real-time rendering while surpassing the past methods in terms of quality.
Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase. But considering both optimization efficiency and expressive capability, we choose to represent dynamic 3D assets by a set of native 4D scene primitives, which is proved to be superior in 4DGS [15].
[…]
4D generation There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [13] deals with a text-to-4D problem by training Hexplane [41] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [14] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, it is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
Recent works have attempted to generate 4D videos from a single input video. For example, TrajectoryCrafter [46] leverages rendered point cloud as the condition of diffusion model to facilitate novel view synthesis. However, the point cloud is projected from the source video and its estimated depth, so there are no points from the back view, limiting the range of novel views. Another work ReCamMaster [47] trains a diffusion model which introduces frame dimension conditioning for input video and camera pose conditioning for target trajectory. But the trajectory is still limited to the front 180. In contrast, our goal is to generate the complete 360 views for a dynamic object.
@@ 3 Preliminary
The 4D Gaussian splatting (4DGS [15]) builds upon the 3D Gaussian splatting technique introduced in [16], originally designed for static scene representation. To address the complexities of dynamic scenes, 4DGS represents each Gaussian defined as:
@@
[…]
@@
Here, $i$ indexes the visible Gaussians sorted by depth, $d$ refers to the direction of the pixel under the view $\mathcal{I}$, and $f_{i}(u,v,t)$ denotes the influence of a Gaussian on this position. To obtain the influence, unlike 3D Gaussian Splatting [16] which directly projects 3D Gaussians to image space, we need condition 4D Gaussians on time and then project them. More specifically, $f(u,v,t)$ is expressed by:
@@
The projection operation $P$ [16, 48] projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by ${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
[…]
Score distillation sampling Score distillation sampling (SDS) is first introduced by DreamFusion [1], which is used to distill the knowledge from a pretrained diffusion model $\epsilon_{\phi}$. Specifically, given an image $I$ rendered from a scene representation (e.g. 3DGS) parameterized by $\theta$, the gradient of SDS loss is calculated as:
@@ 4 Method
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. The input single-view video can be either provided by user or generated by a video generation model. The latter approach extends the application of our method beyond video-to-4D. For example, we can also achieve image-to-4D through a image-to-video diffusion model [17]. As illustrated in Figure 2, it comprises two key components:
@@ 4.1 Image synthesis across views and timestamps
Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Figure 2(a)). Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency. Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Figure 2(c)). This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [49, 8], offering comprehensive information for modeling a dynamic object.
To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video and proceed to generate the remaining views. Our task involves generating multi-view consistent images from a single image for each row. Existing image-to-3D methods, such as SyncDreamer [8], can be leveraged for this purpose. However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images. To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
@@
To further improve temporal resolution, video frame interpolation (e.g., RIFE [50]) can be applied after generating the image matrix $\mathcal{D}$. The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames. This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
@@
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Figure 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.50.50.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio just above 0.50.50.5, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Also note that the results in Figure 3 are not sensitive to the choice.
@@
Aiming for 3D dynamic object modeling, discrete images not suffice. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we have formulated the 4DGS representation model in Section 3, departing from previous slow-to-train 4D reconstruction models [40].
[…]
Optimization In the training of 4DGS, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original loss function, as presented in [15], involves both RGB loss and SSIM loss with balancing weights fixed. Specifically, the loss formulation is defined as:
@@
where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [51], and $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter.
@@
$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}}),$ (12)
@@
The confidence design can guarantee temporal consistency, but achieving high quality of novel view synthesis still remains challenging due to the sparsity of the generated images. Therefore, we also incorporate SDS loss with a relatively small weight $\lambda_{\text{SDS}}$ for smooth transition across different supervised views. We use the image-to-3D diffusion model [17] in SDS loss conditioned on the frame of input video at each timestamp. The total loss function is expressed by:

### v5

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. An intuitive approach is to extend previous image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling. However, this approach would be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model. Importantly, our method can achieve real-time rendering under continuous camera trajectories. To enable robust reconstruction under sparse views, we introduce inconsistency-aware confidence-weighted loss design, along with a lightly weighted score distillation loss. Extensive experiments on both synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D. Our code and video results are publicly available at https://fudan-zvg.github.io/Efficient4D/.
journal: International Journal of Computer Vision
@@ 1 Introduction
Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time. Despite the significant advancement of 3D object generation, existing works Poole et al. (2023); Lin et al. (2023); Chen et al. (2023); Wang et al. (2023) mostly consider static scenes or objects. With the availability of large-scale 3D datasets Deitke et al. (2023b, a), training generalizable models capable of directly generating multi-view images becomes possible Liu et al. (2023b, 2024); Long et al. (2023). These generated images can be turned into a 3D object through reconstruction techniques Mildenhall et al. (2021); Wang et al. (2021). By further augmenting these generated static objects with predefined animations 24, dynamic 3D content can be generated. However, this approach is still limited due to the need for fine-grained meshes as well as rigid restrictions.
@@
Directly generating 4D object/scene content from text description has been recently attempted Singer et al. (2023b). To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation Mildenhall et al. (2021) via score distillation sampling Poole et al. (2023) and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work Jiang et al. (2024) proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as Singer et al. (2023b) this method is also slow to train (e.g., 120 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
@@
To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D. In the first stage, we generate spacetime consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer Liu et al. (2024)) in tandem with frame interpolation. In the second stage, we use these training data to optimize a 4D Gaussian splatting model Yang et al. (2024a). This is an extension of the 3D Gaussian splatting Kerbl et al. (2023), originally designed for static 3D scene representation, with the temporal dimension introduced additionally, allowing for real-time rendering under continuous camera trajectories. Using Gaussian representation brings about further computational efficiency gain, when compared with NeRF based designs (Figure 1). To tackle the challenging discontinuity between the generated sparse frames, we design an inconsistency-aware loss function based on whether there is confidence in the consistency of a frame with its adjacent frames. Along with a lightly weighted score distillation sampling loss for smooth viewpoint transitions, this technique enables robust 4D reconstruction. Notably, although our method takes a single-view video as input, Efficient4D can be easily extended to image-to-4D task by leveraging a image-to-video diffusion model Blattmann et al. (2023).
[…]
Input Consistent4D Jiang et al. (2024): 120 mins Efficient4D (Ours): 10 mins
Input 4DGen Yin et al. (2023): 130 mins Efficient4D (Ours): 10 mins
Input STAG4D Zeng et al. (2024): 70 mins Efficient4D (Ours): 10 mins
@@ 2 Related work
3D generation
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion Poole et al. (2023), introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks Lin et al. (2023); Metzer et al. (2023); Seo et al. (2024); Shi et al. (2023); Chen et al. (2023); Li et al. (2023a); Tsalicoglou et al. (2024); Huang et al. (2023); Wu et al. (2024b); Wang et al. (2023); Zhu and Zhuang (2024) and image-to-3D tasks Xu et al. (2023); Tang et al. (2023); Melas-Kyriazi et al. (2023); Qian et al. (2024) across various dimensions. However, SDS-based approaches often suffer from difficulty to converge and extended optimization times. Conversely, some efficient methods Nichol et al. (2022); Jun and Nichol (2023); Liu et al. (2023b, a, 2024) have emerged. Notably, Point-E Nichol et al. (2022) and Shap-E Jun and Nichol (2023) train models to directly generate 3D point clouds or meshes. Zero123 Liu et al. (2023b) focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or Liu et al. (2023a). Importantly, SyncDreamer Liu et al. (2024) produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
4D representation
Efforts to synthesize videos with free-viewpoint control in dynamic scenes have a well-documented history Zitnick et al. (2004). For example, pre-NeRF Mildenhall et al. (2021) approaches challenges in reconstructing intricate scene details. Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF Pumarola et al. (2021), DeVRF Liu et al. (2022) and HyperNeRF Park et al. (2021), which decouple geometry and motion, utilizing a canonical space and a learned deformation field. DynIBaR Li et al. (2023b) deploys an image-based rendering paradigm for representing long videos with complex camera and object motions. Another group of methods Fridovich-Keil et al. (2023); Cao and Johnson (2023); Shao et al. (2023) adopt tensor decomposition of 4D volumes to represent dynamic 3D scenes.
Recently, Gaussian Splatting Kerbl et al. (2023) has received widespread attention for its real-time high-fidelity rendering, especially its explicit point-based representation which holds great potential in modeling dynamic scenes. Consequently, a significant amount of work has been proposed to explore its extension to dynamic scene modeling. Among them, Deformable 3D Gaussians Yang et al. (2024b) and 4DGaussian Wu et al. (2024a) integrated the deformation field with 3D Gaussian Splatting for the joint learning of the scene’s geometry and dynamics. SC-GS Huang et al. (2024) represents the motion with a set of sparse control points to achieve reconstruction and motion editing. Unlike the previous radiance field-based representations often involving complex training schedules or suffering from slow convergence, these methods can be optimized efficiently, achieving real-time rendering while surpassing the past methods in terms of quality.
Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase. But considering both optimization efficiency and expressive capability, we choose to represent dynamic 3D assets by a set of native 4D scene primitives, which is proved to be superior in 4DGS Yang et al. (2024a).
@@
Recent works have attempted to generate 4D videos from a single input video. For example, TrajectoryCrafter YU et al. (2025) leverages rendered point cloud as the condition of diffusion model to facilitate novel view synthesis. However, the point cloud is projected from the source video and its estimated depth, so there are no points from the back view, limiting the range of novel views. Another work ReCamMaster Bai et al. (2025) trains a diffusion model which introduces frame dimension conditioning for input video and camera pose conditioning for target trajectory. But the trajectory is still limited to the front 180. In contrast, our goal is to generate the complete 360 views for a dynamic object.
@@ 3 Preliminary
The 4D Gaussian splatting (4DGS Yang et al. (2024a)) builds upon the 3D Gaussian splatting technique introduced in Kerbl et al. (2023), originally designed for static scene representation. To address the complexities of dynamic scenes, 4DGS represents each Gaussian defined as:
@@
For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH). Given an arbitrary view $\mathcal{I}$ defined by intrinsic and extrinsic parameters, we render the pixel at position $(u,v)$ with timestamp $t$ by blending visible Gaussians:
@@
Here, $i$ indexes the visible Gaussians sorted by depth, $d$ refers to the direction of the pixel under the view $\mathcal{I}$, and $f_{i}(u,v,t)$ denotes the influence of a Gaussian on this position. To obtain the influence, unlike 3D Gaussian Splatting Kerbl et al. (2023) which directly projects 3D Gaussians to image space, we need condition 4D Gaussians on time and then project them. More specifically, $f(u,v,t)$ is expressed by:
@@
The projection operation $P$ Kerbl et al. (2023); Zwicker et al. (2001) projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by
${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ (7)
where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
Score distillation sampling Score distillation sampling (SDS) is first introduced by DreamFusion Poole et al. (2023), which is used to distill the knowledge from a pretrained diffusion model $\epsilon_{\phi}$. Specifically, given an image $I$ rendered from a scene representation (e.g. 3DGS) parameterized by $\theta$, the gradient of SDS loss is calculated as:
@@ 4 Method
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. The input single-view video can be either provided by user or generated by a video generation model. The latter approach extends the application of our method beyond video-to-4D. For example, we can also achieve image-to-4D through a image-to-video diffusion model Blattmann et al. (2023). As illustrated in Figure 2, it comprises two key components:
@@ 4.1 Image synthesis across views and timestamps
Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Figure 2(a)). Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency. Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Figure 2(c)). This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation Singer et al. (2023a); Liu et al. (2024), offering comprehensive information for modeling a dynamic object.
To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video and proceed to generate the remaining views. Our task involves generating multi-view consistent images from a single image for each row. Existing image-to-3D methods, such as SyncDreamer Liu et al. (2024), can be leveraged for this purpose. However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images. To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
@@
To further improve temporal resolution, video frame interpolation (e.g., RIFE Huang et al. (2022)) can be applied after generating the image matrix $\mathcal{D}$. The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames. This results in a total of M=4K3M=4K-3 images in a column of $\mathcal{D}$.
@@
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Figure 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches $0.5$, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio just above $0.5$, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Also note that the results in Figure 3 are not sensitive to the choice.
@@
Aiming for 3D dynamic object modeling, discrete images not suffice. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we have formulated the 4DGS representation model in Section 3, departing from previous slow-to-train 4D reconstruction models Fridovich-Keil et al. (2023).
Optimization In the training of 4DGS, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original loss function, as presented in Yang et al. (2024a), involves both RGB loss and SSIM loss with balancing weights fixed. Specifically, the loss formulation is defined as:
@@
where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss Wang et al. (2004), and $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter.
@@
[…]
@@
The confidence design can guarantee temporal consistency, but achieving high quality of novel view synthesis still remains challenging due to the sparsity of the generated images. Therefore, we also incorporate SDS loss with a relatively small weight $\lambda_{\text{SDS}}$ for smooth transition across different supervised views. We use the image-to-3D diffusion model Blattmann et al. (2023) in SDS loss conditioned on the frame of input video at each timestamp. The total loss function is expressed by: